OpenStudy (anonymous):

Find the derivative:

OpenStudy (anonymous):

\[\frac{ d }{ dx}\int\limits_{2}^{x}\ln(t^2+1)dt\]

OpenStudy (anonymous):

i need some help to work through it!

OpenStudy (unklerhaukus):

\[\large\frac{\mathrm d}{\mathrm dx}\int\limits_{h(x)}^{g(x}f(t)\mathrm dt=f\big(g(x)\big)\cdot g'(x)-f\big(h(x)\big)\cdot h'(x)\]

OpenStudy (unklerhaukus):

*g(x)

OpenStudy (anonymous):

is that the general rule?

OpenStudy (unklerhaukus):

I think it is called the fundamental theorm of calculus or somesuchnonsense

OpenStudy (anonymous):

my book uses the second fundamental theorem for this chapter haha

OpenStudy (unklerhaukus):

This ol'thing?\[\frac{\mathrm d}{\mathrm dx}\int\limits_{h(x)}^{g(x)}f(x,t)\mathrm dt=f\big(g(x)\big)\cdot g'(x)-f\big(h(x)\big)\cdot h'(x)+\int\limits_{h(x)}^{g(x)}\frac{\partial f(x,t)}{\partial x}\mathrm dt\]

OpenStudy (anonymous):

its this , the second fundamental theorem for constructing antiderivatives \[F(x)=\int\limits_{a}^{x}f(t)dt\]

OpenStudy (anonymous):

the derivative of the integral is the "integrand"

OpenStudy (anonymous):

\[\frac{ d }{ dx}\int\limits_{2}^{x}\ln(t^2+1)dt\] replace \(t\) by \(x\) and get \[\frac{ d }{ dx}\int\limits_{2}^{x}\ln(t^2+1)dt=\ln(x^2+1)\] that is all

OpenStudy (anonymous):

@satellite73 wow that is all? It looks kind of intimidating at first to tell you the truth but this seems pretty easy

OpenStudy (anonymous):

yes, it is pretty easy, the question is only checking if you know that the derivative of the integral is the integrand that is all

OpenStudy (anonymous):

sounds good